Find all solutions to the equation.
cos2x + 2 cos x + 1 = 0

Answer by stanbon(75887)

Find all solutions to the equation.
cos2x + 2 cos x + 1 = 0
(2cos^2-1) + 2cos + 1 = 0
2cos^2 + 2cos = 0
(cos)(2cos+2) = 0
cos(x) = 0 or cos(x) = -1
-----
x = pi/2 or (3/2)pi or x = pi
--------------------------------
Cheers,
Stan H.
================